OpenStudy (anonymous):

What is the value of x? (1 point) 5 2.5 7.5 10

OpenStudy (stamp):

2.5

OpenStudy (stamp):

x/(2x+5)=(x-2)/(2x-1) x(2x-1)=(x-2)(2x+5) 2xx-x=2xx+x-10 2x=10 x=5